A man managing a photocopying establishment for $48,000 per year decides to open his own duplicating...

  1. A man managing a photocopying establishment for $48,000 per year decides to open his own duplicating shop. His revenue during the first year of operations is $190,000 and his expenses are as follows:

Salaries to hired help                                                   $50,000

Suplies                                                                              $30,000

Rent                                                                                  $25,000

Utilites                                                                             $3,000

Interest on bank loan                                                  $25,000                                                                           Calculate (a) the explicit costs (b) the implicit costs, (c) the business profit, (d) economic profit and (e) the normal return on investment in this business.

For Problem 3, see p. 14. Note that the normal return or normal profit is an amount paid to the owner that is just necessary to get the owner to be in business; i.e., just necessary to cover the owner’s implicit cost.
Answer #1

(a)

Explicit cost ($) = Salaries + Supplies + Rent + Utilities + Interest on bank loan = 50,000 + 30,000 + 25,000 + 3,000 + 25,000

= 133,000

(b)

Implicit cost ($) = Income given up from managing photocopy shop = 48,000

(c)

Business profit ($) = Revenue - Explicit cost = 190,000 - 133,000 = 57,000

(d)

Economic profit ($) = Business profit - Implicit cost = 57,000 - 48,000 = 9,000

(e)

Normal return ($) = Implicit cost = 48,000